OpenStudy (anonymous):

find the solution to the system of equations by the addition method. -5x-7y=-57 -0.9x + 0.7y= -11.2 0.1x + 0.7y = -4.2

OpenStudy (anonymous):

ignore the -5x part

OpenStudy (saifoo.khan):

-0.9x + 0.7y= -11.2 0.1x + 0.7y = -4.2 - - + ============== -1.0x + 0 = -7 -x = -7 x = 7

OpenStudy (saifoo.khan):

Now, substitute it in any equation, you will get the value for y